OpenStudy (anonymous):

The area of a regular octagon is 35 cm. what is the area of a regular octagon with sides three times as large?

TheSmartOne (thesmartone):

Welcome to OpenStudy! :D @ariana_keri Well if we multiply the side by 3. Then that means that we will be multiplying the total area by \(\sf 3^2=9\) So your answer will be \(\sf 35\times 9=?\) (If it wanted volume, then we would have to multiply it by \(\sf 3^3=27\))

OpenStudy (anonymous):

thank you soo much @TheSmartOne

TheSmartOne (thesmartone):

Np :) Also since you are new here, check out this tutorial for new OpenStudiers so you know how to get around the site :) http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8